OpenStudy (anonymous):

Find the rectangular coordinates for the point whose polar coordinates are given. (6√2, 5π/3)

OpenStudy (anonymous):

\[x=rcos(\theta) , y=rsin(\theta)\]

OpenStudy (anonymous):

\[ \sin(5\pi/3) = \sin(\pi/3) = (\frac{\sqrt{3}}{2})\]

OpenStudy (anonymous):

For (6sqrt2))cos(5pi/3) I get 4.242640607. Do I round that up or what else?

OpenStudy (anonymous):

what form do i put it in

OpenStudy (anonymous):

\[\cos(5\pi/3) = - \cos(\pi/3) = -\frac{1}{2}\]

OpenStudy (anonymous):

you can get exact values!,

OpenStudy (anonymous):

they want exact values

OpenStudy (anonymous):

do i not put the 6sqrt2 in the front

OpenStudy (anonymous):

yeah, exact values

OpenStudy (anonymous):

no, I was just finding the angles

OpenStudy (anonymous):

the answers u gave are wrong.

OpenStudy (anonymous):

is 1/2 the x value? and sqrt3/2 y

OpenStudy (anonymous):

\[ x =-6\sqrt{2} \frac{1}{2} , y= 6\sqrt{2}\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

lol plz dont tell u you put them in :(

OpenStudy (anonymous):

^ thats the FINAL answer

OpenStudy (anonymous):

i did.

OpenStudy (anonymous):

well simplify it a bit

OpenStudy (anonymous):

x = -3sqrt(2) , y= 3sqrt(6)

OpenStudy (anonymous):

the answer for x seems very obscure

OpenStudy (anonymous):

the other one's r wrong to

OpenStudy (anonymous):

no their not

OpenStudy (anonymous):

that's what the my hw program says. i hate it :P

OpenStudy (amistre64):

x = 3sqrt(2) y = -3sqrt(6)

OpenStudy (anonymous):

no,

OpenStudy (amistre64):

5[pi/3 is in the 4th quadrant; +x, -y

OpenStudy (anonymous):

wait lols , maybe

OpenStudy (amistre64):

360-60 :)

OpenStudy (anonymous):

yeh I was thinking 2pi/3 :|

OpenStudy (anonymous):

amistre to the rescue: Thank you for all ur help eng :)

OpenStudy (amistre64):

elec did al the hard stuff; got rid of the errors I would have done lol